Solve the following equations:
(3/2) l = 2/3.........

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

First we have to multiply both the side by 2,
= (3l/2) × 2 = (2/3) × 2
= 3l = (4/3)
Now,
We have to divide both the side by 3,
Then we get,
= 3l/3 = (4/3)/3
= l = (4/3) × (1/3)
= x = (4/9)

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve this question as follows :
3/2 I = 2/3
3 I / 2 = 2/3
3 I = 4/3
I = 4/9
 
Hope it helps!
Thank you ✌️